games, where 31 of them were won by the home team. Use a 0.10 significance level to test H:p-0.5 H,: p 0.5 H:p<0.5 dentify the test stafistic for this hypothesis test The tost statistic for this hypothesis test is Round to two decimal places as needed.) The P-value for this hypotess test is (Round to three decimal places as needed.) O A. Fail to reject Ho- There is not suficient evidenoe to support the claim that the probability of thes home leam winning is greater than one-hat OB. Fail to reyect Ho. There s steent evidence to support the aim that the probably ofthe home team wenngisgreater ran one-hat. OC. Reject H There is sufficient evidence to support the clsim thait the probabilty of the home team wiening s greater han one-half. OD, Rayed Ke There i e, sMient evide ce to support he damthat the probablity ofthe home team wnnng is geater Pan onehalf. FS F7 esc F2 F4
Answer #1

Ans:

Option B is correct.

H0: p=0.5

H1:p>0.5

sample proportion=31/53=0.5849

Test statistic:

z=(0.5849-0.5)/sqrt(0.5*(1-0.5)/53)

z=1.24

p-value=P(z>1.24)=0.107

As,p-value>0.1,we fail to reject H0.

Option A is correct.